A line k passes through the points (4,2) and (-4,4). A second line h passes through the points (-8,1) and (8,-3). Is line k parallel to line h? Why or Why not?
Answer by checkley77(12844)

Y=mX+b
(4,2) and (-4,4).
m=(Y2-Y1)/(X2-X1)
m=(4-2)/(-4-4)
m=2/-8
m=-1/4 ans. for this slope.
(-8,1) and (8,-3).
m=(-3-1)/(8+8)=-4/16=-1/4 ans. for this slope.
These lines are parallel because the slopes are equal.